# Deep Learning for Subspace Regression

### v2

@@ Abstract
It is often possible to perform reduced order modelling by specifying linear subspace which accurately captures the dynamics of the system. This approach becomes especially appealing when linear subspace explicitly depends on parameters of the problem. A practical way to apply such a scheme is to compute subspaces for a selected set of parameters in the computationally demanding offline stage and in the online stage approximate subspace for unknown parameters by interpolation. For realistic problems the space of parameters is high dimensional, which renders classical interpolation strategies infeasible or unreliable. We propose to relax the interpolation problem to regression, introduce several loss functions suitable for subspace data, and use a neural network as an approximation to high-dimensional target function. To further simplify a learning problem we introduce redundancy: in place of predicting subspace of a given dimension we predict larger subspace. We show theoretically that this strategy decreases the complexity of the mapping for elliptic eigenproblems with constant coefficients and makes the mapping smoother for general smooth function on the Grassmann manifold. Empirical results also show that accuracy significantly improves when larger-than-needed subspaces are predicted. With the set of numerical illustrations we demonstrate that subspace regression can be useful for a range of tasks including parametric eigenproblems, deflation techniques, relaxation methods, optimal control and solution of parametric partial differential equations.
@@ 1 Introduction
Code used for training, evaluation and dataset generation is available at https://github.com/VLSF/subreg.
@@ 2.1 Definition of subspace regression problem
Let $V(r):\mathbb{R}^{p}\rightarrow\text{Gr}(k,n)$ be a function that maps the space of parameters $r\in\mathbb{R}^{p}$ to the subset of grassmanian represented as the set of tall full rank matrices $V(r)$. We assume that parameters $r$ are sampled from distribution $r\sim p_{r}$ and that dataset $\mathcal{D}=\left\{\left(r_{1},V_{1}\right),\dots,\left(r_{m},V_{m}\right)\right\}$ of $m$ i.i.d. samples is available. For a given parametric model $W_{\theta}:\mathbb{R}^{p}\rightarrow\text{Gr}(r,n),r\geq k$ we want to identify parameters $\theta^{\star}$ such that $W_{\theta^{\star}}(r)$ approximates $V(r)$. We formulate this task as optimisation problem
$\theta^{\star}=\arg\min_{\theta}\left(\underset{r\sim p_{r}}{\mathbb{E}}\left[L\left(W_{\theta}(r),V(r)\right)\right]\right)\simeq\arg\min_{\theta}\left(\frac{1}{m}\sum_{i=1}^{m}L\left(W_{\theta}(r_{i}),V_{i}\right)\right).$ (1)
@@
Aside from unusual invariance requirement (2), optimisation problem (1) is a standard machine learning formulation of regression problems which can be solved with stochastic optimisation for arbitrary model $W_{\theta}(r)$ that admits efficient evaluation of gradients.
@@ 2.2 Examples of subspace regression problem
$\Delta\psi(x)+U(x)=E\psi(x),\,\left\|\psi\right\|_{2}<\infty.$ (3)
One way to find eigenpairs is to approximates $\psi(x)$ by a finite series $\psi(x)=\sum_{i=1}^{K}\alpha_{i}\phi_{i}(x)$ and enforce Petrov-Galerkin condition that residual is orthogonal to all $\phi_{i}(x)$. Continuous problem (3) reduces to eigenproblem for Hermitian matrix and can be solved in $O\left(K^{3}\right)$ operations [32]. For eigenproblems we are typically interested only in extremal eigenspaces corresponding to either smallest or largest eigenvalues. In this case it is desirable to select a small number of basis functions $\phi_{i}(x)$ that approximate sufficiently well the subspace of interest. When eigenproblem (3) is solver repeatedly for many potential functions $U(x)$ this lead us to subspace regression problem (1) used to approximate the mapping $U(x)\rightarrow\left\{\sum_{i=1}^{K}\alpha_{i}\phi_{i}(x)=f,\alpha_{i}\in\mathbb{R}\right\}$. [^1] Suitable discretisation of parametrisation of both $\phi$ and $U$ is assumed. When this mapping is learned from a set of examples, eigenproblems for unobserved potentials $U$ can be solved efficiently, since low-dimensional candidate subspace for eigenfunctions is available.
@@
When equation (5) is discretised with finite difference or finite element method, it reduces to linear problem $A\phi=f$, where AA is large sparse matrix and $\phi,f$ are discretised solution and right-hand side of (5). To exploit sparsity of AA one can solve linear equation with relaxation method. General relaxation method split matrix additively $A=D+C$, where $D$ is regular with known inverse [27]. Given the split, if iteration scheme $x^{n+1}=x^{n}+D^{-1}\left(b-Ax^{n}\right)$ is convergent, steady state is exact solution $x=A^{-1}b$. Convergence is linear and its rate is defined by spectral radius of error propagation matrix $I-D^{-1}A$. To improve convergence rate, one can augment iterative method with coarse-grid correction [33]. This techniques allows one to remove influence of leading subspace $V$ of $I-D^{-1}A$ by solving small reduced linear system for error equation $V^{\top}AVe=r$, where $e$ and $r$ are error and residual in the subspace $\mathcal{S}(V)$. Naturally, subspace regression (1) for this problem approximates the mapping $A\rightarrow\mathcal{S}(V)$ or $k(x)\rightarrow\mathcal{S}(V)$ for linear systems resulting from equation (5).
Deflation for conjugate gradient. Krylov subspace methods provide a more systematic way to solve large sparse linear systems [27]. For linear system with symmetric positive definite matrix AA resulting from discretisation of equation (5), the method of choice is conjugate gradient (CG) [15]. Similar to other Krylov methods, on step $r$, CG identify optimal solution within Krylov subspace $\mathcal{K}_{r}=\text{span}\left\{b,Ab,\dots,A^{r-1}b\right\}$, where span refers to the subspace formed by linear combinations of vectors in the set. Since powers of AA are involved, the most readily available vectors are from the subspaces with large eigenvalues [28]. To improve convergence of method it is reasonably to include eigenspaces $V$ with small eigenvalues to the approximation space $\mathcal{K}_{r}$. The resulting method is deflated CG and the approximation space is $\mathcal{K}_{r}\cup\mathcal{S}(V)$ [29]. In analogy with previous example, the subspace regression problem considered approximates $A\rightarrow\mathcal{S}(V)$ or $k(x)\rightarrow\mathcal{S}(V)$, but this time $V$ spans eigenspaces with small eigenvalues of matrix AA.
@@ 3.1 Loss functions
    Loss function $L_{1}(A,B)=p-\left\|Q_{B}^{\top}Q_{A}\right\|_{F}^{2}$ satisfies requirements (2), where $A=Q_{A}R_{A},B=Q_{B}R_{B}$ are QR decompositions of AA and $B$, $\left\|\cdot\right\|_{F}$ is Frobenius norm.
@@
    Let $z\in\mathbb{R}^{k}$ be a random variable with zero mean and identity covariance matrix. Loss functions $L_{2}(A,B;z)=\min_{u}\left\|Au-Q_{B}z\right\|_{2}^{2}$ does not depend on the choice of AA from $\lceil A\rceil$, where $B=Q_{B}R_{B}$ is QR decomposition.
@@
Loss $L_{1}$ is essentially the same as the difference of orthogonal projectors. Loss $L_{2}$ introduces two modifications: (i) projector in Riemannian distance is replaced with error of least squares problem; (ii) to remove second projector, stochastic Hutchinson trace estimation is used. Reformulation with least square problem allows one to use normal equation, and various tools from randomised numerical linear algebra, e.g., randomised preconditioned Cholesky-QR [14], blendenpik solver [2], and sketching [35]. We will see in Section 4, that loss function $L_{2}(A,B)$ based on normal equation scales better than $L_{1}(A,B)$ with the increase of subspace size.
@@ 3.2 Subspace embedding
In the definition of subspace regression problem (1) we allowed to approximate function $\mathbb{R}^{p}\rightarrow\text{Gr}(k,n)$ by function $\mathbb{R}^{p}\rightarrow\text{Gr}(r,n)$ where $r\geq k$. We call this strategy subspace embedding. It is justified because of two unique properties of regression and interpolation on grassmanian: (i) inclusion of vector subspaces is well-defined; (ii) after interpolation subspace is used to construct a reduced model. From the latter property one may expect similar or improved accuracy when the predicted subspace from $\text{Gr}(r,n)$ contains target subspace from $\text{Gr}(k,n)$.
We will show empirically in Section 4 that subspace embedding significantly improves accuracy and generalisation gap. Here we argue that prediction of larger-than-needed subspaces align well with inductive bias of neural networks known as f-principle or spectral bias [36],. F-principle is an observation that neural networks tend to learn smoothed versions of the target functions. As we show below, embedding techniques may improve smoothness of learned function.
@@
Let $V(t):\mathbb{R}\rightarrow\text{Gr}(k,n)$ be continuously differentiable on $t\in[0,T]$, $V(t)^{\top}V(t)=I_{k}$. It is always possible to construct piecewise continuous function $W(t):\mathbb{R}\rightarrow\text{Gr}(r,n)$, $r>k$, $W(t)^{\top}W(t)=I_{r}$ such that $\frac{1}{2}\left\|W(t)W(t)^{\top}-V(t)V(t)^{\top}\right\|_{F}^{2}-\frac{r-k}{2}$ is arbitrary small and $\left\|\dot{W}(t)\right\|_{F}^{2}\leq\left\|\dot{V}(t)\right\|_{F}^{2}$, where inequality is strict for all points where $\left\|\dot{V}(t)\right\|_{F}^{2}\neq 0$.
@@ 3.3 Complexity of parametric eigenproblem
Suppose eigenvectors of (7) are ordered according to the increase of eigenvalues. Let $\phi_{k}$ be an eigenvector on position $k$, let $V_{k}$ be an eigenspace spanned by vectors on positions up to $k$. Consider mappings $F:a_{1},\dots,a_{D}\rightarrow\phi_{k}$ and $G:a_{1},\dots,a_{D}\rightarrow V_{k}$.
@@
    $F,G$ are piecewise constant functions that take $_{F}(k,D),\#_{G}(k,D)<\infty$ distinct values.
@@
    Let $W_{l},l>1$ be a subspace obtained by union of $V_{l}$ for distinct $a_{1},\dots,a_{D}$. Number of vectors in $W_{l}$ is $_{F}(l,D)+1$.
@@
    $_{F}(k,D)\sim\frac{1}{(D-1)!}k\left(\log k\right)^{D-1}$ for fixed $D$ and large $k$ .
@@
    $_{G}(k,D)\geq\frac{1}{(D-1)!}k^{D-1}$ for fixed $D$ and large $k$.
@@
    $_{G}(k,D)\geq\frac{1}{(k-1)!}D^{k-1}$ for fixed $k$ and large $D$.
@@
Theorem 3 suggest that for problem (7) mappings from coefficient to $k$-th eigenvector $\phi_{k}$ or subspace $V_{k}$ are piecewise constant functions with rapidly growing number of constant regions when either $k$ or $D$ increases. The complexity (the number of regions) of the subspace prediction problem exceeds the complexity of $k$-th eigenvector prediction. However, results also suggest that the number of unique eigenvectors within $V_{k}$ grows at the same rate as the number of eigenvectors on position $k$. This means, the large number of distinct regions in $G$ comes from a large number of possible combinations of an asymptotically small number of vectors. Given that, the complexity of mapping from coefficients to subspace decreases, if one predicts subspace of larger dimension $\widetilde{V}_{k}\supseteq V_{k}$. For example, if one is willing to predict subspace of dimension $_{F}(k,D)$ in place of $V_{k}$ of dimension $k$, the mapping $a_{1},\dots,a_{D}\rightarrow\widetilde{V}_{k}\supseteq V_{k}$ may be chosen to have constant value.
@@ 4 Numerical experiments
We present several numerical experiments to corroborate our theoretical findings. The discussion of control problems appears in Appendix I.

### v3

# arXiv:2509.23249v3[cs.LG] 27 Feb 2026

@@ Abstract
It is often possible to perform reduced order modelling by specifying linear subspace which accurately captures the dynamics of the system. This approach becomes especially appealing when linear subspace explicitly depends on parameters of the problem. A practical way to apply such a scheme is to compute subspaces for a selected set of parameters in the computationally demanding offline stage and in the online stage approximate subspace for unknown parameters by interpolation. For realistic problems the space of parameters is high dimensional, which renders classical interpolation strategies infeasible or unreliable. We propose to relax the interpolation problem to regression, introduce several loss functions suitable for subspace data, and use a neural network as an approximation to high-dimensional target function. To further simplify a learning problem we introduce redundancy: in place of predicting subspace of a given dimension we predict larger subspace. We show theoretically that this strategy decreases the complexity of the mapping for elliptic eigenproblems with constant coefficients and makes the mapping smoother for general smooth function on the Grassmann manifold. Empirical results also show that accuracy significantly improves when larger-than-required subspaces are predicted. With the set of numerical illustrations we demonstrate that subspace regression can be useful for a range of tasks including parametric eigenproblems, deflation techniques, relaxation methods, optimal control and solution of parametric partial differential equations.
@@ 2.1 Definition of subspace regression problem
Let $V:\mathbb{R}^{p}\rightarrow\text{Gr}(k,n)$ be a function that maps the space of parameters $r\in\mathbb{R}^{p}$ to the subset of grassmanian represented as the set of tall full rank matrices $V(r)$. We assume that parameters $r$ are sampled from distribution $r\sim p_{r}$ and that dataset $\mathcal{D}=\left\{\left(r_{1},V_{1}\right),\dots,\left(r_{m},V_{m}\right)\right\}$ of $m$ i.i.d. samples is available. For a given parametric model $Y_{\theta}:\mathbb{R}^{p}\rightarrow\text{Gr}(r,n),r\geq k$ we want to identify parameters $\theta^{\star}$ such that $Y_{\theta^{\star}}(r)$ approximates V(r)V(r) [^1] Note, that we allow $Y_{\theta}$ to have more columns than target $V$. In this context approximation is understood in terms of subspace inclusion $\mathcal{S}\left(V\right)\subset\mathcal{S}\left(W_{\theta}\right)$. As we explain later, redundancy introduced this way can significantly improve accuracy.. We formulate this task as optimisation problem
$\theta^{\star}=\arg\min_{\theta}\left(\underset{r\sim p_{r}}{\mathbb{E}}\left[L\left(Y_{\theta}(r),V(r)\right)\right]\right)\simeq\arg\min_{\theta}\left(\frac{1}{m}\sum_{i=1}^{m}L\left(Y_{\theta}(r_{i}),V_{i}\right)\right).$ (1)
@@
Aside from unusual invariance requirement (2), optimisation problem (1) is a standard machine learning formulation of regression problems which can be solved with stochastic optimisation for arbitrary model $Y_{\theta}(r)$ that admits efficient evaluation of gradients.
@@ 2.2 Examples of subspace regression problem
$\Delta\psi(x)+U(x)\psi(x)=E\psi(x),\,\left\|\psi\right\|_{2}<\infty,$ (3)
where $U(x)$ is potential energy and $E$ is energy of the system. One way to find eigenpairs is to approximate $\psi(x)$ by a finite series $\psi(x)=\sum_{i=1}^{K}\alpha_{i}\phi_{i}(x)$ and enforce Petrov-Galerkin condition that residual is orthogonal to all $\phi_{i}(x)$. Continuous problem (3) reduces to eigenproblem for Hermitian matrix and can be solved in $O\left(K^{3}\right)$ operations [33]. For eigenproblems we are typically interested only in extremal eigenspaces corresponding to either smallest or largest eigenvalues. In this case it is desirable to select a small number of basis functions $\phi_{i}(x)$ that approximate sufficiently well the subspace of interest. When eigenproblem (3) is solved repeatedly for many potential functions $U(x)$ this lead us to subspace regression problem (1) used to approximate the mapping $U(x)\rightarrow\left\{f(x):f(x)=\sum_{i=1}^{K}\alpha_{i}\phi_{i}(x),\alpha_{i}\in\mathbb{C}\right\}$. [^2] Suitable discretisation of parametrisation of both $\phi$ and $U$ is assumed. That is, we wish to predict subspace spanned by first $K$ eigevectors. When this mapping is learned from a set of examples, eigenproblems for unobserved potentials $U$ can be solved efficiently, since low-dimensional candidate subspace for eigenfunctions is available.
@@
When equation (5) is discretised with finite difference or finite element method, it reduces to linear problem $A\phi=f$, where $A$ is large sparse matrix and $\phi,f$ are discretised solution and right-hand side of (5). To exploit sparsity of $A$ one can solve linear equation with relaxation method. General relaxation method split matrix additively $A=D+C$, where $D$ is regular with known inverse [28]. Given the split, if iteration scheme $x^{n+1}=x^{n}+D^{-1}\left(b-Ax^{n}\right)$ is convergent, steady state is exact solution $x=A^{-1}b$. Convergence is linear and its rate is defined by spectral radius of error propagation matrix $I-D^{-1}A$. To improve convergence rate, one can augment iterative method with coarse-grid correction [34]. This techniques allows one to remove influence of leading subspace formed by colums of matrix $V$ of $I-D^{-1}A$ by solving small reduced linear system for error equation $V^{\top}AVe=r$, where $e$ and $r$ are error and residual in the subspace $\mathcal{S}(V)$. Naturally, subspace regression (1) for this problem approximates the mapping $A\rightarrow\mathcal{S}(V)$ or $k(x)\rightarrow\mathcal{S}(V)$ for linear systems resulting from equation (5).
Deflation for conjugate gradient. Krylov subspace methods provide a more systematic way to solve large sparse linear systems [28]. For linear system with symmetric positive definite matrix $A$ resulting from discretisation of equation (5), the method of choice is conjugate gradient (CG) [15]. Similar to other Krylov methods, on step $r$, CG identify optimal solution within Krylov subspace $\mathcal{K}_{r}=\text{span}\left\{b,Ab,\dots,A^{r-1}b\right\}$, where span refers to the subspace formed by linear combinations of vectors in the set. Since powers of $A$ are involved, the most readily available vectors are from the subspaces with large eigenvalues [29]. To improve convergence of method it is reasonably to include eigenspaces $V$ with small eigenvalues to the approximation space $\mathcal{K}_{r}$. The resulting method is deflated CG and the approximation space is $\mathcal{K}_{r}\cup\mathcal{S}(V)$ [30]. In analogy with previous example, the subspace regression problem considered approximates $A\rightarrow\mathcal{S}(V)$ or $k(x)\rightarrow\mathcal{S}(V)$, but this time $V$ spans eigenspaces with small eigenvalues of matrix $A$.
@@ 3.1 Loss functions
    Loss function $L_{1}(A,B)=p-\left\|Q_{B}^{\top}Q_{A}\right\|_{F}^{2}$ satisfies requirements (2), where $A=Q_{A}R_{A},B=Q_{B}R_{B}$ are reduced QR decompositions of $A$ and $B$, $\left\|\cdot\right\|_{F}$ is Frobenius norm [^3] Reduced QR decomposition of tall full rank matrix $A\in\mathbb{R}^{n\times k}$ is a factorisation $A=Q_{A}R_{A}$, where $Q_{A}\in\mathbb{R}^{n\times k}$ has orthonormal columns, $R_{A}\in\mathbb{R}^{k\times k}$ is upper triangular with nonzero elements on the diagonal..
@@
    Let $z\in\mathbb{R}^{k}$ be a random variable with zero mean and identity covariance matrix. Loss functions $L_{2}(A,B;z)=\min_{u}\left\|Au-Q_{B}z\right\|_{2}^{2}$ does not depend on the choice of $A$ from $\lceil A\rceil$, where $B=Q_{B}R_{B}$ is QR decomposition.
@@
Loss $L_{1}$ is essentially the same as the difference of orthogonal projectors. Note, that $L_{1}(A,B)\geq 0$ with equality reached if and only if matrices $A$ and $B$ share the same columns space, since in this case $\left\|Q_{B}^{\top}Q_{A}\right\|_{F}^{2}=p$. Loss $L_{2}$ introduces two modifications: (i) projector in Riemannian distance is replaced with error of least squares problem; (ii) to remove second projector, stochastic Hutchinson trace estimation is used. Reformulation with least square problem allows one to use normal equation, and various tools from randomised numerical linear algebra, e.g., randomised preconditioned Cholesky-QR [14], blendenpik solver [2], and sketching [36]. We will see in Section 4, that loss function $L_{2}(A,B)$ based on normal equation scales better than $L_{1}(A,B)$ with the increase of subspace size.
@@ 3.2 Subspace embedding
In the definition of subspace regression problem (1) we allow to approximate function $\mathbb{R}^{p}\rightarrow\text{Gr}(k,n)$ by function $\mathbb{R}^{p}\rightarrow\text{Gr}(r,n)$ where $r\geq k$. We call this strategy subspace embedding. It is justified because of two unique properties of regression and interpolation on grassmanian: (i) inclusion of vector subspaces is well-defined; (ii) subspace, predicted by regression model or interpolated by standard techniques, is used to construct a reduced model. From the latter property one may expect similar or improved accuracy when the predicted subspace from $\text{Gr}(r,n)$ contains target subspace from $\text{Gr}(k,n)$.
We will show empirically in Section 4 that subspace embedding significantly improves accuracy and generalisation gap. Here we argue that prediction of larger-than-needed subspaces align well with inductive bias of neural networks known as f-principle or spectral bias [37]. F-principle is an observation that neural networks tend to learn smoothed versions of the target functions. As we show below, embedding techniques may improve smoothness of learned function.
@@
Let $V:\mathbb{R}\rightarrow\text{Gr}(k,n)$ be continuously differentiable on $t\in[0,T]$, $V(t)^{\top}V(t)=I_{k}$. It is always possible to construct piecewise continuous function $W:\mathbb{R}\rightarrow\text{Gr}(r,n)$, $r>k$, $W(t)^{\top}W(t)=I_{r}$ such that $\frac{1}{2}\left\|W(t)W(t)^{\top}-V(t)V(t)^{\top}\right\|_{F}^{2}-\frac{r-k}{2}$ is arbitrary small and $\left\|\dot{W}(t)\right\|_{F}^{2}\leq\left\|\dot{V}(t)\right\|_{F}^{2}$, where inequality is strict for all points where $\left\|\dot{V}(t)\right\|_{F}^{2}\neq 0$.
@@ 3.3 Complexity of parametric eigenproblem
Suppose eigenvectors of (7) are ordered according to the increase of eigenvalues. Let $\phi_{k}$ be an eigenvector on position $k$, let $V_{k}$ be an eigenspace spanned by vectors on positions up to $k$. Consider mappings $F_{k}:a_{1},\dots,a_{D}\rightarrow\phi_{k}$ and $G_{k}:a_{1},\dots,a_{D}\rightarrow V_{k}$.
@@
    $F_{k},G_{k}$ are piecewise constant functions that map real numbers to elements of sets $S_{F_{k}},S_{G_{k}}$. Sets $S_{F_{k}},S_{G_{k}}$ are finite with $_{F_{k}}(k,D),\#_{G_{k}}(k,D)$ distinct elements.
@@
    Let $W_{l},l>1$ be a subspace obtained by union of $V_{l}$ for distinct $a_{1},\dots,a_{D}$. Number of vectors in $W_{l}$ is $_{F_{k}}(l,D)+1$.
@@
    $_{F_{k}}(k,D)\sim\frac{1}{(D-1)!}k\left(\log k\right)^{D-1}$ for fixed $D$ and large $k$ .
@@
    $_{G_{k}}(k,D)\geq\frac{1}{(D-1)!}k^{D-1}$ for fixed $D$ and large $k$.
@@
    $_{G_{k}}(k,D)\geq\frac{1}{(k-1)!}D^{k-1}$ for fixed $k$ and large $D$.
@@
Theorem 3 suggests that for problem (7) mappings from coefficient to $k$-th eigenvector $\phi_{k}$ or subspace $V_{k}$ are piecewise constant functions with rapidly growing number of constant regions when either $k$ or $D$ increases. The complexity (the number of regions) of the subspace prediction problem exceeds the complexity of $k$-th eigenvector prediction. However, results also suggest that the number of unique eigenvectors within $V_{k}$ grows at the same rate as the number of eigenvectors on position $k$. This means, the large number of distinct regions in $G$ comes from a large number of possible combinations of an asymptotically small number of vectors. Given that, the complexity of mapping from coefficients to subspace decreases, if one predicts subspace of larger dimension $\widetilde{V}_{k}\supseteq V_{k}$. For example, if one is willing to predict subspace of dimension $_{F}(k,D)$ in place of $V_{k}$ of dimension $k$, the mapping $a_{1},\dots,a_{D}\rightarrow\widetilde{V}_{k}\supseteq V_{k}$ may be chosen to have constant value.
@@ 4 Numerical experiments
We present several numerical experiments to corroborate our theoretical findings. The discussion of control problems appears in Appendix I. All numerical results are reported as single-run metrics without explicit error bars. To study sensitivity to train-test split we perform several dedicated experiments. Results suggest that variance is low and does not affect main conclusions. The details are available in Appendix J. For most problems we report relative error measured in percents. For predicted quantity $\widetilde{v}$ and ground truth value $v$ it reads $100\%\times\left\|v-\widetilde{v}\right\|_{2}\big/\left\|v\right\|_{2}$. For eigenvalue problems the numerator is $\mathbb{Z}_{2}$ adjusted as explained below. For iterative methods we report convergence plots for relative error, without additional factor $100\%$.